A SEARCH FOR INTERSTELLAR CARBON-CHAIN ALCOHOL HC_{4}OH IN THE STAR FORMING RE-GION L1527

\underline{MITSUNORI ARAKI}, Department of Chemistry, Faculty of Science Division I, Tokyo University of Science, 1-3 Kagurazaka, Shinjuku-ku, Tokyo, 162-8601, Japan; SHURO TAKANO, Nobeyama Radio Observatory, 462-2 Nobeyama, Minamimaki, Minamisaku, Nagano, 384-1305, Japan; HIROMICHI YAMABE NAOHIRO KOSHIKAWA, KOICHI TSUKIYAMA, Department of Chemistry, Faculty of Science Division I, Tokyo University of Science, 1-3 Kagurazaka, Shinjuku-ku, Tokyo, 162-8601, Japan; AYA NAKANE, TOSHIAKI OKABATYASHI, Department of Chemistry, Faculty of Science, Shizuoka University, 836 Oya, Suruga-ku, Shizuoka 422-8529, Japan; ARISA KUNIMATSU and NOBUHIKO KUZE, Department of Materials and Life Sciences, Faculty of Science and Technology, Sophia University, 7-1 Kioi-cho, Chiyoda-ku, Tokyo, 102-8554, Japan.

We have made a sensitive search for the rotational transitions of carbon-chain alcohol HC_{4}OH with the frequency ragne from 21.2 to 46.7 GHz in the star forming region L1527 in Taurus with rich carbon-chain chemistry. The incentive of this observation was a laboratory detection of HC_{4}OH by the microwave spectroscopy. Despite achieving an rms of several mK in antenna temperature by the 45m telescope at Nobeyama Radio Observatory, the searche for HC_{4}OH was negative, leading to a 5 sigma upper limit corresponding to the column density of 4 × 10^{12} cm^{−2} based on the excitation temperature of 12.3 K. The upper limit indicates that the [HC_{4}-OH]/[HC_{4}-CN] ratio is less than 1.0. The ratio suggests that the cyanide species with carbon-chain structure is dominant in comparison with the hydroxyl one in L1527, which can be the opposite case of saturated compounds, e.g. CH_{3}OH and CH_{3}CN, in hot cores and dark clouds.